OpenStudy (lollylau):

Series of trial-and-error questions: Find all the unknowns for: 1. x^y=y^x 2. x^y=y^x+1 where all the unknowns are integers.

OpenStudy (lollylau):

?????????!!!!!!!!

OpenStudy (anonymous):

i guess \[y=x\] might work for first one

OpenStudy (lollylau):

could be...

OpenStudy (lollylau):

let's say they aren't, some actual answers please...

OpenStudy (anonymous):

what is wrong with \[3^3=3^3\]?

OpenStudy (anonymous):

2 and 4 works for the first one.

OpenStudy (lollylau):

Nothing! Pratu got it right :) (Well you too...)

OpenStudy (mr.math):

Sorry, I thought it was a system of two equations.

OpenStudy (lollylau):

oh.

OpenStudy (anonymous):

ooooooooooooooh. i was wondering...

OpenStudy (mr.math):

I agree with satellite on the first one, the solution of the first one is all integers (x,y) such that x=y. For the second one, we have (x,y)=(2,1) is a solution.

OpenStudy (mr.math):

(3,2) is another solution.

OpenStudy (anonymous):

pratu had \[2^4=4^2\] so that works as well. my guess is there are no others, but that is a guess

OpenStudy (lollylau):

There are no others.

OpenStudy (lollylau):

no proof tho. just know that for SURE :)

OpenStudy (lollylau):

Mr. Math, you're missing one more answer. (which i know)

OpenStudy (lollylau):

(solution)

OpenStudy (lollylau):

ans. in half a min.

OpenStudy (mr.math):

(1,0).

OpenStudy (lollylau):

YAY!

OpenStudy (mr.math):

Actually any solution in the form (x,0), x is a positive integer.

OpenStudy (lollylau):

Can we do this for +2, +3, +4 etc.

OpenStudy (mr.math):

Yes.